A meterstick of negligible mass has a 0.20-kg mass
Answer #1

About the fulcrum the net torque on the meter stick will be zero

So let say the fulcrum is at distance "x" from the origin

Torque of left mass = torque of right mass

0.20g*(x - 35) = 0.40g*(75 - x)

(x - 35) =2*(75 - x)

3x =150 + 35 = 185

x = 61.67 cm